Front Cover: The upper panels present observations of the prototypical merger remnant, NGC 7252. The optical light is rendered in magenta while the neutral hydrogen gas is rendered in blue. The observational material is fully described in Chapter III. At the bottom of the page is Hubble’s famous tuning-fork diagram [Hubble, E., 1936, “The Realm of the Nebulae” (New Haven, Yale Univ. Press).], after a left–right inversion. This figure, which captures the spirit of the merger hypothesis, was presented by François Schweizer at the Heidelberg interacting galaxy conference [Schweizer, F., 1990, “Dynamics and Interactions of Galaxies”, R. Wielen editor, p. 60 (Heidelberg, Springer-Verlag)].

Back Cover: N-body calculation of the NGC 7252 encounter, as viewed from the plane of the sky, from 64 h^{−1} Myr before orbital periapse to 580 h^{−1} Myr after, which is found to be the best match to the neutral hydrogen data. See Chapter IV for details.

The Fate of Gas in Merging Disk Galaxies

John Eugene Hibbard

Submitted in partial fulfillment of the requirements for the degree of Doctor of Philosophy

in the Graduate School of Arts and Sciences

COLUMBIA UNIVERSITY

ABSTRACT

The Fate of Gas in Merging Disk Galaxies

John Eugene Hibbard

In this thesis I present a multi-wavelength study of five merging disk systems. I directly test the hypothesis that the merger of two atomic gas rich spirals will leave an atomic gas poor elliptical, by measuring the quantities, distribution, and kinematics of the atomic gas along an evolutionary disk-disk merger sequence. These data are supplemented by deep broad band R and narrow band Hα imaging to completely map the faint stellar distribution and regions of current star formation. I identify several trends along the sequence, the most noteworthy of which are the presence of plumes of ionized gas emanating from the minor-axis in the early stages of merging, and a very low central atomic gas content in the late stage mergers. Overall, large quantities of both gas and starlight are sent to distant radii. Since this material evolves on very long time scales, it may leave observable signatures of the merger for many billions of years. Substructures are found within the expanding tails, with properties which suggested that they might evolve into dwarf-like systems. A complete analysis of one of the systems, the prototypical merger remnant NGC 7252, is presented. A comparison of the gas kinematics and optical and gaseous morphologies suggests that much tidal material remains bound to the system and is presently streaming back into the remnant body. This hypothesis is supported by detailed numerical simulations of the encounter, which successfully reproduce the H i velocity data. The observations of atomic gas streaming back into the remnant while the remnant body remains atomic gas free, coupled with the fact that this system has photometric properties characteristic of ellipticals, leads us to conclude that some mergers are likely to evolve into elliptical galaxies. The simulation is followed beyond the best fit time to study the future evolution of the merger remnant, and we find that half of the remaining tidal material will fall back towards the remnant within the next few Gigayears. Neutral hydrogen observations should therefore be an effective method for uncovering aged merger remnants. Finally I explore some differences between the archetypical disk-disk mergers and the most luminous infrared galaxies, which are believed to be ongoing mergers. The observations may indicate necessary prerequisites for triggering luminous infrared activity in mergers, such as specific orbital geometries or Hubble types for the progenitors.

TABLE OF CONTENTS

Contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . i Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ix

Chapter I: Thesis Objectives . . . . . . . . . . . . . . . . . . . . . . . . . . 3

1 Motivation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2 Historical Perspectives . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.1 Bridge and Tail Building . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 Observational Validation . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.2.1 Definition of a Merger Remnant . . . . . . . . . . . . . . . . . . . . . 11 2.2.2 Shells, Fine Structures, and Aging Merger Remnants . . . . . . . . . . . . 12 2.2.3 Ultraluminous IRAS Galaxies and Mergers . . . . . . . . . . . . . . . . 15 3 Formulation of The Thesis . . . . . . . . . . . . . . . . . . . . . . . . . 18 4 Layout of The Thesis . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

Chapter II: H i, H ii, and R-band Observations of a Galactic Merger Sequence …. 27

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 2 The Evolutionary Sequence . . . . . . . . . . . . . . . . . . . . . . . . . 29 3 Observations and Data Reduction . . . . . . . . . . . . . . . . . . . . . . 32 3.1 Radio Observations . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 3.2 Optical Observations . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 4 Results . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.1 Global Results . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 4.2 Individual Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 4.2.1 Arp 295 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 4.2.2 NGC 4676 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.2.3 NGC 520 & UGC 957 . . . . . . . . . . . . . . . . . . . . . . . . . 51 4.2.4 NGC 3921 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.2.5 NGC 7252 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 5 Observed Trends Along the Sequence . . . . . . . . . . . . . . . . . . . . 56 5.1 Plume of Ionized Hydrogen . . . . . . . . . . . . . . . . . . . . . . . . 57 5.2 Optical Morphology . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 5.3 Central Atomic Gas Content . . . . . . . . . . . . . . . . . . . . . . . 58 5.4 Extent of Tidal Features . . . . . . . . . . . . . . . . . . . . . . . . . 59 6 Discussion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 6.1 Effects of Merging on the Inner Regions . . . . . . . . . . . . . . . . . . 62 6.1.1 Starlight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 6.1.2 Ionized Hydrogen . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 6.1.3 Cold Gas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 6.2 Effects of Merging on the Outer Regions . . . . . . . . . . . . . . . . . . 65 6.2.1 Returning Tidal Material . . . . . . . . . . . . . . . . . . . . . . . . 65 6.2.2 Clumps in Tails . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 6.2.3 Differences Between Gas and Light Distributions . . . . . . . . . . . . . . 67 6.3 Future Evolution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 6.3.1 Evolution of Tails . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 6.3.2 Evolution of Starlight . . . . . . . . . . . . . . . . . . . . . . . . . . 70 6.3.3 Evolution of Gas Content . . . . . . . . . . . . . . . . . . . . . . . . 71 6.3.4 Other Fates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 7 Summary and Conclusions . . . . . . . . . . . . . . . . . . . . . . . . . 72

Chapter III: Cold, Warm, and Hot Gas in the Late-Stage Merger NGC 7252 . . . . . 81

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 2 Observations & Data Reduction . . . . . . . . . . . . . . . . . . . . . . . 84 2.1 Radio Observations . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 2.2 Optical Observations . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 2.3 X-ray Observations . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 3 Results . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 3.1 Global H i Properties . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 3.1.1 Spatial Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . 88 3.1.2 Kinematics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 3.2 Gas in the Main Body . . . . . . . . . . . . . . . . . . . . . . . . . . 96 3.2.1 Spatial Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . 96 3.2.2 Kinematics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 3.3 Gas & Star Formation in the Tidal Tails . . . . . . . . . . . . . . . . . . 107 3.3.1 Large Scale Properties of Stars & Gas . . . . . . . . . . . . . . . . . . 107 3.3.2 Giant H ii Regions . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 4 Discussion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 4.1 Gas Distribution and Progenitors . . . . . . . . . . . . . . . . . . . . . 114 4.2 Kinematic Structure of the Tidal Features . . . . . . . . . . . . . . . . . 115 4.3 Dearth of Central Atomic Gas . . . . . . . . . . . . . . . . . . . . . . . 116 4.4 From Spiral to Elliptical? . . . . . . . . . . . . . . . . . . . . . . . . . 117 4.5 Clumps in Tidal Tails . . . . . . . . . . . . . . . . . . . . . . . . . . 119 5 Summary and Conclusions . . . . . . . . . . . . . . . . . . . . . . . . . 120

Chapter IV: Dynamical Modeling of N7252 and the Return of Tidal Material …. 127

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 2 Numerical Technique . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 3 Simulation fit to the HI Data . . . . . . . . . . . . . . . . . . . . . . . . 132 4 Properties of the Simulation Model . . . . . . . . . . . . . . . . . . . . . 138 4.1 Time Evolution and Orbital Structure of the Tails . . . . . . . . . . . . . . 138 4.2 The Future Evolution of NGC 7252 . . . . . . . . . . . . . . . . . . . . 142 5 Discussion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 5.1 Lack of Neutral Hydrogen in NGC 7252 . . . . . . . . . . . . . . . . . . 146 5.2 Return of Tidal Material . . . . . . . . . . . . . . . . . . . . . . . . . 149 6 Conclusions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150

Chapter V: Disk-Disk Mergers and the Luminous IR Galaxies . . . . . . . . . . . 155

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 2 Comparison Between IR Galaxies and Disk-Disk Mergers . . . . . . . . . . . . 157 2.1 The Comparison Samples . . . . . . . . . . . . . . . . . . . . . . . . . 157 2.2 The Comparison . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 3 Discussion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 4 Future Work . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

Chapter VI: Concluding Remarks . . . . . . . . . . . . . . . . . . . . . . . 175

1 The Fate of Gas during Disk-Disk Mergers . . . . . . . . . . . . . . . . . . 175 2 Answers to the Posed Questions . . . . . . . . . . . . . . . . . . . . . . . 176 3 Remaining Questions . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 4 Final Comment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186

Appendices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 189

I VLA Observations . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 II Details on the Arp 295 System . . . . . . . . . . . . . . . . . . . . . . . 197 III Details on the NGC 4676 System . . . . . . . . . . . . . . . . . . . . . 205 IV Details on the NGC 520 System . . . . . . . . . . . . . . . . . . . . . . 213 V Details on the NGC 3921 System . . . . . . . . . . . . . . . . . . . . . 223 VI Details on Other Systems . . . . . . . . . . . . . . . . . . . . . . . . 232 VII Clumps in Tidal Tails . . . . . . . . . . . . . . . . . . . . . . . . . . 235 VIII Details of the Modeling Effort . . . . . . . . . . . . . . . . . . . . . . 243

LIST OF FIGURES AND TABLES

Chapter I

Fig. I-1. Simulation of Arp 295 (Toomre & Toomre 1972) . . . . . . . . . . . . . 5 Fig. I-2. Simulation of NGC 4676 (Toomre & Toomre 1972) . . . . . . . . . . . . 5 Fig. I-3. Location of Tidal Material in Original Disk (Toomre & Toomre 1972) …. 7 Fig. I-4. The Toomre Sequence (Toomre 1977) . . . . . . . . . . . . . . . . . . 8 Fig. I-5. Spectral light and Kinematics of NGC 7252 (Schweizer 1982) . . . . . . . 12 Fig. I-6. ULIR Galaxies (Joseph & Wright 1985). . . . . . . . . . . . . . . . . 14 Fig. I-7. The ULIR Merging Timeline (Casoli et al. 1991) . . . . . . . . . . . . . 16 Fig. I-8. Alternate SFR Timelines (Mihos & Hernquist 1994) . . . . . . . . . . . 17

Chapter II

Fig. II-1. The Merging Sequence . . . . . . . . . . . . . . . . . . . . . . . 31 Fig. II-2. Deep R-band Images of the Merging Sequence . . . . . . . . . . . . . 38 Fig. II-3. H i, H ii, and R-band Images of the Merging Sequence (color) . . . . . Plate 1 Fig. II-4. H i Greyscales and R-band Contours of the Merging Sequence . . . . . . 39 Fig. II-5. Intensity-weighted velocity fields of the Merging Sequence (color) … Plate 2 Fig. II-6. H ii and R-band Images of the Main Bodies . . . . . . . . . . . . . . 40 Fig. II-7.a. H i Mass-to-Light Ratios for Early-Stage Tidal Features . . . . . . . . 48 Fig. II-7.b. H i Mass-to-Light Ratios for Late-Stage Tidal Features . . . . . . . . 49 Fig. II-8. Fraction Central Gas and Light . . . . . . . . . . . . . . . . . . . 59 Fig. II-9. Subtended Area . . . . . . . . . . . . . . . . . . . . . . . . . . 61

Table II-1. VLA Observing Parameters . . . . . . . . . . . . . . . . . . . . 33 Table II-2. CTIO/KPNO Observing Parameters . . . . . . . . . . . . . . . . 34 Table II-3. Global Properties of Systems . . . . . . . . . . . . . . . . . . . . 42 Table II-4 Properties of Inner Regions . . . . . . . . . . . . . . . . . . . . . 44 Table II-5. Properties of Tidal Features . . . . . . . . . . . . . . . . . . . . 46

Chapter III.

Fig. III-1. H i, H ii, and Stellar Distribution in NGC 7252 . . . . . . . . . . . . . 89 Fig. III-2. H i and R-band Image of Companion to NGC 7252 . . . . . . . . . . . 90 Fig. III-3. H i C+D Array Channel Maps of NGC 7252 . . . . . . . . . . . . . . 93 Fig. III-4. H i Velocity Structure . . . . . . . . . . . . . . . . . . . . . . . 95 Fig. III-5.a. X-ray Distribution on R-band Light . . . . . . . . . . . . . . . . 97 Fig. III-5.b. H ii Distribution on R-band Light . . . . . . . . . . . . . . . . . 98 Fig. III-5.c. H i Distribution on Model Subtracted R-band Light . . . . . . . . . . 99 Fig. III-5.d. H i Distribution on R-R Colors . . . . . . . . . . . . . . . . . . . 100 Fig. III-5.e. X-ray, H ii, and H i Distribution on R-band Light . . . . . . . . . . . 102 Fig. III-5.f. ^{12}CO(1→0), H ii, and H i Distribution on R-band Light . . . . . . . . . 103 Fig. III-6. R-band Radial Brightness Profile . . . . . . . . . . . . . . . . . . 104 Fig. III-7. H i and H ii Velocities on B-R Colors . . . . . . . . . . . . . . . . . 106 Fig. III-8. Details of NW Tail . . . . . . . . . . . . . . . . . . . . . . . . 110 Fig. III-9. Details of E Tail . . . . . . . . . . . . . . . . . . . . . . . . . . 111

Table III-1. VLA Observing Parameters of NGC 7252 . . . . . . . . . . . . . . 85 Table III-2. CTIO/KPNO Observing Parameters of NGC 7252 . . . . . . . . . . 86 Table III-3. ROSAT Observing Parameters/Spectral Fit for NGC 7252 . . . . . . 87 Table III-4. Global Properties of NGC 7252 . . . . . . . . . . . . . . . . . . 91 Table III-5. Colors in Main Body of NGC 7252 . . . . . . . . . . . . . . . . 101 Table III-6. Properties Along Tails of NGC 7252 . . . . . . . . . . . . . . . . 109 Table III-7. Properties of Tail Clumps of NGC 7252 . . . . . . . . . . . . . . . 112

Chapter IV.

Fig. IV-1. H i Velocity Structure . . . . . . . . . . . . . . . . . . . . . . . 133 Fig. IV-2. Simulation Match to HI Data . . . . . . . . . . . . . . . . . . . . 135 Fig. IV-3. Simulation Match to Optical Image . . . . . . . . . . . . . . . . . 137 Fig. IV-4. Time Evolution of Model . . . . . . . . . . . . . . . . . . . . . . 139 Fig. IV-5. Orbital Structure of Tails . . . . . . . . . . . . . . . . . . . . . . 141 Fig. IV-6. Development of NW Tail (color) . . . . . . . . . . . . . . . . . Plate 3 Fig. IV-7. Long Term Time Evolution of Tidal Tails . . . . . . . . . . . . . . . 143 Fig. IV-8. Future Evolution of Present Day Tails (color) . . . . . . . . . . . Plate 4 Fig. IV-9. Original Distribution of Tail Material . . . . . . . . . . . . . . . . 147

Table IV-1. Timescales and Quantities of Returning Tidal Material . . . . . . . . 145

Chapter V.

Fig. V-1. Images of ULIR Sample . . . . . . . . . . . . . . . . . . . . . . . 156 Fig. V-2. Images of LIR Sample . . . . . . . . . . . . . . . . . . . . . . . 158 Fig. V-3. Images of Toomre Sequence . . . . . . . . . . . . . . . . . . . . . 160 Fig. V-4. Illustration of ULIR-merger Timeline . . . . . . . . . . . . . . . . . 165 Fig. V-5. Simulated R-band Image of the Merger Sequence at z=0.1 . . . . . . . . 167

Table V-1. IR Selected Samples . . . . . . . . . . . . . . . . . . . . . . . . 162 Table V-2. Optically Selected Sample . . . . . . . . . . . . . . . . . . . . . 163 Table V-3. Intercomparison of Samples . . . . . . . . . . . . . . . . . . . . 164

Appendices

Table A-1. Properties of H i Detected Systems . . . . . . . . . . . . . . . . . 231

Appendix I

Fig. AI-1. VLA Configuration and UV Sampling Pattern . . . . . . . . . . . . . 192 Fig. AI-2. Beams for Different Weighting Functions . . . . . . . . . . . . . . . 194

Appendix II

Fig. AII-1. R-band, H i, and H ii emission from Arp 295 . . . . . . . . . . . . . 196 Fig. AII-2. Detail on Southern Tail of Arp 295a . . . . . . . . . . . . . . . . . 199 Fig. AII-3. H i Channel Maps of Arp 295 . . . . . . . . . . . . . . . . . . . . 200 Fig. AII-4. H i Channel Maps of Northern Companions to Arp 295 . . . . . . . . 202 Fig. AII-5. H i Intensity and Intensity-Weighted Velocity Field of Arp 295 . . . . . 203 Fig. AII-6. B − R Map of Arp 295 Region (color) . . . . . . . . . . . . . . . Plate 5

Appendix III

Fig. AIII-1. R-band, H i, and H ii emission from NGC 4676 . . . . . . . . . . . . 204 Fig. AIII-2. Detail on Northern Tail of NGC 4676a . . . . . . . . . . . . . . . 207 Fig. AIII-3. H i Channel Maps of NGC 4676 . . . . . . . . . . . . . . . . . . 208 Fig. AIII-4. H i Intensity and Intensity-Weighted Velocity Field of N4676 . . . . . . 210

Appendix IV

Fig. AIV-1. R-band, H i, and H ii emission from NGC 520 . . . . . . . . . . . . 212 Fig. AIV-2. Detail on Inner Region of NGC 520 . . . . . . . . . . . . . . . . 215 Fig. AIV-3. H i Channel Maps of NGC 520 System . . . . . . . . . . . . . . . 216 Fig. AIV-4. H i Channel Maps of UGC 957 Region . . . . . . . . . . . . . . . 218 Fig. AIV-5. H i Intensity and Intensity-Weighted Velocity Field near U957 . . . . . 219 Fig. AIV-6. H i Intensity and Intensity-Weighted Velocity Field of N520 . . . . . . 220 Fig. AIV-7. B − R Map of NGC 520 (color) . . . . . . . . . . . . . . . . . Plate 6

Appendix V

Fig. AV-1. R-band, H i, and H ii emission from NGC 3921 . . . . . . . . . . . . 222 Fig. AV-2. Detail on Inner Region of NGC 3921 . . . . . . . . . . . . . . . . 225 Fig. AV-3. H i Channel Maps of NGC 3921 System . . . . . . . . . . . . . . . 226 Fig. AV-4. R-band and H i emission from NGC 3916 . . . . . . . . . . . . . . . 228 Fig. AV-5. H i Intensity and Intensity-Weighted Velocity Field of N3921 . . . . . . 229 Fig. AV-6. Position-Velocity Profile for NGC 3921 . . . . . . . . . . . . . . . 230

Appendix VI

Fig. AVI-1. Photometric Properties of NGC 7252 . . . . . . . . . . . . . . . . 233 Fig. AVI-2. B − R Map of NGC 7252 (color) . . . . . . . . . . . . . . . . . Plate 7 Fig. AVI-3. B − R Map of NGC 4038/9 (color) . . . . . . . . . . . . . . . . Plate 8

Appendix VII

Fig. AVII-1. Clumps in NGC 7252 . . . . . . . . . . . . . . . . . . . . . . 238 Fig. AVII-2. Clumps in NGC 3921 . . . . . . . . . . . . . . . . . . . . . . 240 Fig. AVII-3. Clumps in NGC N520 . . . . . . . . . . . . . . . . . . . . . . 241

Table AVII-1. Measured Properties of Clumps . . . . . . . . . . . . . . . . . 236

Appendix VIII

Table A-2. Restricted 3-Body Simulation Parameter Search . . . . . . . . . . . 246 Table A-3. N-Body Simulation Parameter Search . . . . . . . . . . . . . . . . 247

Fig. AVIII-1. Starting Geometry . . . . . . . . . . . . . . . . . . . . . . . 256 Fig. AVIII-2. Time Evolution from Above Orbital Plane . . . . . . . . . . . . . 257 Fig. AVIII-3. Accretion Rate of Tidal Material . . . . . . . . . . . . . . . . . 258 Fig. AVIII-4. Color-coded NW disk at Periapse (color) . . . . . . . . . . . . Plate 9 Fig. AVIII-5. Color-coded E − V_{x} Plane for Fig. 4 (color) . . . . . . . . . . . Plate 10 Fig. AVIII-6. Evolution of Ring of Particles in Roche Frame . . . . . . . . . . . 259 Fig. AVIII-7. Velocity Vectors in Tails . . . . . . . . . . . . . . . . . . . . . 260 Fig. AVIII-8. E, J, Period, R_{peri} Time Evolution . . . . . . . . . . . . . . . . 261 Fig. AVIII-9. Period-R_{peri} Plane for E and NW Tails . . . . . . . . . . . . . . 262 Fig. AVIII-10. Sky View of NGC 7252 at T=72 and T=172 . . . . . . . . . . . 263 Fig. AVIII-11. Column Density Evolution . . . . . . . . . . . . . . . . . . . 264 Fig. AVIII-12. Orbital Plane Evolution, t=0.7–5.8 Gyr . . . . . . . . . . . . . 266 Fig. AVIII-13. (r, v_{r}) Plane Evolution, t=0.7–5.8 Gyr . . . . . . . . . . . . . . 267 Fig. AVIII-14. Orbital Plane Evolution, t=3.3–13.5 Gyr . . . . . . . . . . . . . 268 Fig. AVIII-15. Orbital Plane Evolution, t=3.3–13.5 Gyr, Expanded View . . . . . 269 Fig. AVIII-16. Appearance of Tidal Features at t=3.26 Gyr . . . . . . . . . . . 270 Fig. AVIII-17. Appearance of Tidal Features at t=13.5 Gyr . . . . . . . . . . . 271 viii

ACKNOWLEDGEMENTS

My graduate career got off to a false start in 1985 when, in search of possible Masters thesis topics, I was told that “interacting galaxies were not that important”, and to work on something else. Partly due to this response, I spun out of academia and landed in L.A., devising ways to kill men in tanks. I owe my eternal gratitude to my advisor, Jacqueline van Gorkom, for having the generousity and faith to offer me the opportunity to return to my first astronomical love, interacting galaxies. She also indulged my every whim as I defined this thesis, and supported me shamelessly in its pursuit, providing me direction when she could, and, when she couldn’t, providing me with the contacts and collaborators who could. Most of all I thank Jacqueline for sharing her love of science and discovery with me time and again, and for serving as a constant reminder of why I am in the field at all. I also thank her for enjoying a good argument, and the ability not to take it personally. I am indebted to several individuals for getting me started in this whole game. First off, to Steven Shore, for introducing me to Ring Galaxies in my senior year at CWRU, and for teaching me that to be a scientist means to never stop asking questions. I thank NRAO for two summer student appointments early in my graduate career, and the staff and scientists of the VLA for making that time so special and exciting. My thanks go out again to Jacqueline who, during one of those appointments, offered me my first H i spectral line data set to work on. The object was Mkn 348, and it should be no surprise that since then I have been hooked on 21cm spectral line mapping of extragalactic objects. I owe Mike Rich my sincere thanks for sending me to Heidelberg to attend the most influential conference of my career, on the Dynamics and Interactions of Galaxies, at which my thesis was first outlined, and for much support throughout my graduate career at Columbia. I would also like to thank the participants of that conference for proving undeniably that interacting galaxies are that important. Mostly, I thank François Schweizer for his many invaluable contributions over the years. It was François who, at the Heidelberg conference, told Jacqueline and I about the single dish H i detection of NGC 7252 by Dupraz et al. (1990), and who excitedly talked us into proposing a follow-up VLA observation, netting the results presented on the cover of this thesis and detailed in Chapter III. Many times since that day, François’ enthusiasm infected me as he took me on his personal tour of the astronomical zoo as documented on the original plate material in his office, and it was he who suggested the objects of study for this thesis. I thank François for the ten-hour discussions on mergers and remnants and ripples and dynamics; for greatly improving my thesis proposal, my papers, and this thesis by being a “nit-picking swiss”; and, like Jacqueline, for reminding me that there are scientist who are deeply devoted to their work purely for the joy of discovery and the love of knowing. After Jacqueline and François, the people I have learned most from during these years are Puragra (Raja) Guhathakurta and J. Christopher Mihos. I thank Raja for sharing his art of communication with me, for sticking to his convictions when I was a complete pain in the ass, and for having the patience to wait for me to see the error of my ways. I hope I have learned half of what he has tried to teach me. I thank Chris for being the only theoretician I met who had the guts to send me ACTUAL DATA following my inquiries, and for being as interested as I was in the questions I asked; an attitude which eventually led to his collaboration on Chapter IV of this thesis. I was given scientific and personal support from many fronts over these past six years. The staff and visitors of the VLA must be especially recognized for their indulgence, and interest and support — Ed Fomalont, Miller Goss, Juan Uson, Barry Clark, Michael Rupen, The Carillis (Chris, Adrianna, Michael, and Country Joe), Theresa McBride, John Spargo, Chris Depree, Peggy and Rick Perley, The Wunk, Emilio, Eileen Latasa, the operators, Larry Brothers, Doug Wood, Jay Nowicki, the many summer students, and many many others. I also thank those who shared their scientific expertise with me, to Dave Westpfahl for showing me how REAL astronomers take optical data (answer: sitting next to the telescope in the freezing cold and swinging it around with your bare hands), to Raja again for a wonderful trip to Chili and for sharing his mosaicing techniques, to Remo Tilanus for his help with CO observations which turned up some interesting non-detections which I never had time to write up, and to Lars Hernquist and Josh Barnes for correcting my many misconceptions about theoretical implications. I am particularly grateful to Lars for donating the CPUs and allowing me to come to Santa Cruz to conduct the NGC 7252 simulation. To Nick Scoville and Min Yun for many fascinating discussions on the ultraluminous infrared galaxies, and for a free and generous exchange of ideas about how to tackle this project, which should led to a resolution to many of the questions brought up in Chapter V of this thesis. To Neil deGrasse Tyson for teaching me the value of having every reference at your finger tips, to Rob Olling and Pauline McMahon for helping me tackle the sometimes frustrating world of AIPS and spectral line observing and, as fellow van Gorkom graduate students, for commiseration and moral support when she would give one of us “that look”; to Dave Schiminovich for many discussions on shells, to Hong Sheng Zhao for many discussions B&T, and to Thijs van der Hulst, The Nicest Astronomer and a true pioneer in this field. I am also in debt to and in love with the sometimes kind, sometimes sad, sometimes funny, and always memorable souls at Columbia. To Mother Rob Olling, for taking care of me when I was ill, feeding me when I was hungry, and making the transition to unix as painless as it can be. To Andy Rasmussen and Janna Levin for the food orgies, the space cake, the poetry readings, the fights with flaccid white male <expletive deleted>’s, the New Year Eves, and for a deep and lasting friendship. To Pauline McMahon and Chris Chimichanga (Cimaszewski), Edhead Moran, to the memory of Marta Torres, to Phil Yecko, Cheryl Hartup, Steve Houspian, Misa Chappel, Leigh Fredrickson, my brother David, Tom Hamilton and Bronwen Jones, David Helfand, Ben “yes my buddy” Sternleib, “The Goils”, Drew Heckler, and the Hungarian Cafe, for the beers, the singing, the music, the movies, the coffee, the dancing, the arguments, the happy hours, the East Village treks, and for generally making NYC a great place to be a grad student. I also thank the many folks from places afar for keeping me sane. To Albert Melfo, Mitch Duran and Huss (the Police Dog), Sally Anne Weeker, Tiziana Venturi, Jayanne English, Rich Gelderman, Gabriella Vehar, and Sarah Israelit; to Jeff Hemmer, Dave Hudak, Johnny Pelquer, the Chicken with a Hat, and all the other LA/SC/Mexico grogs; to these and all others who offered me shelter from the storm over these years. I spent the last 6 months that it took to assemble this thesis at the Institute for Astronomy, and I thank the members of that community for many fruitful and enlightening discussions which helped me to fully develop Chapter V of this thesis. To Dave Sanders, Josh Barnes, John Kormendy, and to the participants of the Interacting Galaxy Seminar. Back at Columbia I must thank both Mildred Garcia and JvG for their patience and perseverance, without which I could not have graduated; and a special thanks to Dave Schiminovich, who made the “Dutch Styled” version of this thesis possible. Finally, I extend my sincere thanks to the unfailing patience, faith, and support of the family Hibbard. And with my deepest love and appreciation, I thank my wife Leslie Levin and her baby, my son, Alexsander, for doing what no one ever could by showing me that there are things more important than astronomy and the galaxies I study.

John E. Hibbard April, 1995